# Incorporating Control Inputs in the Estimation of Continuous Mobile Robot Trajectories and Continuum Robot Shapes

### v1

@@ III-B Localization Evaluation
III-C Results
The results are summarized in Tab. I, including the estimation errors and computation times for increasing times between landmark distance measurements . When including a discrete state at every odometry update, both methods achieve similar results. The resulting accuracies differ only by a few mm and less than half of a degree, while the WNOA GP prior baseline achieves shorter computation times. When including discrete states only when landmark distance measurements occur, the proposed prior formulation using inputs significantly outperforms the baseline. While the traditional WNOA GP prior is able to achieve competitive accuracies for short time distances , it fails to accurately represent the local shape of the trajectory when landmark measurements occur less frequently ( s). The proposed GP prior with inputs is able to maintain the estimation accuracy for longer times. Here the error only starts to significantly increase for s. Both methods show significantly reduced computation times when excluding states without landmark distance measurements in the batch estimation. However, the computation time increases when using a prior with inputs as increases. In these cases, more states have to be recovered using the GP interpolation schemes, which requires additional effort for the proposed method to solve the Magnus expansion and integral terms.
We conclude that using the proposed GP prior with inputs presents an interesting use case, as it is able to maintain the estimation accuracy, while significantly reducing the computational effort. This is thanks to its ability to maintain the local shape of the odometry over longer time intervals without needing to include discrete states in the batch estimation. We can exploit this advantage, as long as does not grow too large, in which case the accuracy and computation times are not competitive with the WNOA GP prior baseline, even without excluding discrete states from the batch estimation.
This property is further highlighted in Fig. 4, which shows state estimation results for the first 25% of the mobile robot dataset, while considering landmark distance measurements every s and including discrete estimation states only at these measurement times. The proposed method is able to accurately represent the local shape of the trajectory between the discrete states. In contrast, the baseline, which relies on the constant-velocity prior, cannot accurately capture the local shape of the trajectory and instead produces a highly smoothed result. This leads to considerable inaccuracies in the resulting state estimation.
IV Continuum Robot Shape Estimation
To further validate the proposed GP prior formulation, continuum robot experiments were conducted. As mentioned earlier, the methods above can be directly applied to the quasi-static estimation of a continuum robot’s shape (see [5]).
IV-A Experimental Setup
For the experiments, we are using the tendon-driven continuum robot (TDCR) prototype and sensor setup presented in [5] (see Fig. 5, left). For sensing, we consider a single six-degree-of-freedom pose measurement at the robot’s tip using an electromagnetic tracking sensor (Aurora v3, Northern Digital Inc., Canada). Ground-truth is obtained using a coordinate measurement arm with an attached laser probe (FARO Edge with FARO Laser Line Probe HD, FARO Technologies Inc., USA), from which the discrete pose of each spacer disk can be reconstructed. Six simple bending configurations of the continuum robot were recorded to calibrate the sensor and ground-truth frames. Using the same procedure outlined in [5], the remaining calibration error at the tip of the robot results in mm.
IV-B Evaluation
For evaluation, we consider nine different TDCR configurations. Three sets of actuation values, defined as applied tendon tensions, are considered within this set, meaning that each set is shared by three configurations. In each configuration, an additional unknown force is applied to the robot’s tip to further deflect it.
For state estimation, we now wish to compare the conventional prior formulation from [5] with the newly proposed one considering inputs. For continuum robots, there are two possibilities to incorporate known inputs. One can either consider known strains (equivalent to known velocities) or known distributed actuation loads (equivalent to accelerations, when additionally considering the robot material’s stiffness matrix), either of which could be provided by another open-loop continuum robot model.
We use the second approach and convert the known applied tendon tensions to moments acting on the robot backbone, considering the simplified point-moment model discussed in [24]. Using this model and the known stiffness of the robot’s Nitinol backbone (using properties reported by the manufacturer), we compute discrete moments applied to the locations at which the tendons terminate. We then approximate these discrete moments with piecewise-linear inputs. While this simplified model exhibits inaccuracies in out-of-plane bending scenarios [24], we deem it sufficiently accurate to be fused with additional measurements. Alternatively, more complex actuation load distributions could be incorporated, as long as they can be expressed in the robot’s body frame independently from the state variables.
Both the traditional prior from [5] as well as the newly proposed prior considering inputs are then fused with a single pose measurement at the tip of the robot. As an additional degenerate case, we further consider the scenario in which only the robot tip’s position can be measured. For every case, we compute the RMSE and maximum errors for the robot’s position and orientation along its shape, considering the discrete ground-truth poses at each disk.
IV-C Results
@@
The results for each of the nine configurations, as well as their average, are reported in Tab. II. In each case, the state estimation problem could be solved in 5-10 ms. Fig. 5, middle and right, additionally showcases the results for the first configuration, when considering a pose measurement. It can be seen that due to the ability to retain the local shape of the prior, our proposed method outperforms the conventional GP formulation in every case, with two exceptions. In these cases, relatively high external forces were applied to the continuum robot, which resulted in shapes that significantly differ from the utilized prior, making it less helpful.
V Conclusion
@@
In this work, we propose an extension to the WNOA GP prior used in continuous-time batch state estimation. Our new formulation allows us to incorporate both known velocity and acceleration inputs, as long as they can be expressed in the robot’s body frame independently of the robot state. We specifically present the case of piecewise-linear inputs, for which closed-form approximations can be obtained to represent motions in . Experiments showcase the use cases of the method for both mobile robot trajectory and continuum robot shape estimation. It is shown that considering inputs in the GP prior is particularly helpful to maintain the local shape of the trajectory between discrete state observations, specifically when the motion differs from the constant-velocity assumption of the traditional WNOA GP prior.
Future work will focus on further evaluating the approach on mobile robots with more aggressive motions, such as present in drone racing, in which cases the traditional WNOA GP prior is less accurate. Further, it will be investigated whether inputs can also be considered in the WNOJ GP motion prior as an alternative.

### v2

@@ 4.2 Evaluation
4.3 Results
@@
The results for each of the nine configurations, as well as their average, are reported in Tab. 2. In each case, the state estimation problem could be solved in 5-10 ms. Fig. 5, middle and right, additionally showcases the state estimation results for the first configuration, both when considering a full pose measurement and a position-only measurement of the continuum robot’s tip. It can be seen that due to the ability to retain the local shape of the prior, our proposed method outperforms the conventional GP formulation in every case, with two exceptions. In these cases, relatively high external forces were applied to the continuum robot, which resulted in shapes that significantly differ from the utilized prior, making it less helpful. Nevertheless, the results in these cases remain largely comparable to those obtained by the baseline method.
In the remaining configurations, the accuracy of the proposed method is significantly higher than that of the baseline method, especially in cases where only the position of the robot’s tip is measured. In these instances, the baseline method’s shape estimates largely deviate from the ground-truth, with worst-case maximum errors exceeding 3 cm and 45∘. This further underscores the proposed method’s effectiveness in situations with sparse sensing, as the additional inputs contribute to a more informed prior, resulting in more a accurate state estimation. This is also apparent in the example shown in Fig. 5.
5 Conclusion
@@
In this work, we propose an extension to the WNOA GP prior used in continuous-time batch state estimation. Our new formulation allows us to incorporate both known velocity and acceleration inputs, as long as they can be expressed in the robot’s body frame independently of the robot state. We specifically present the case of piecewise-linear inputs, for which closed-form approximations can be obtained to represent motions in . Experiments showcase the use cases of the method for both mobile robot trajectory and continuum robot shape estimation. It is shown that considering inputs in the GP prior is particularly helpful to maintain the local shape of the trajectory between discrete state observations, specifically when the motion differs from the constant-velocity assumption of the traditional WNOA GP prior. Future work will focus on further evaluating the approach on mobile robots with more aggressive motions, such as present in drone racing, in which cases the traditional WNOA GP prior is less accurate. Further, it will be investigated whether inputs can also be considered in the WNOJ GP motion prior as an alternative.
@@ Acknowledgements
Author Contribution
SL and TDB both conceptualized the research question, method and experiments. SL wrote the majority of the paper, derived and implemented the proposed method and carried out the qualitative and quantitative evaluations. TDB is the senior author on this paper. He regularly participated in discussions on the proposed research, method and results, and edited the paper.
Financial Support
This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC).
Competing Interests Declaration
The authors declare none.